### v1

# Active Visual Thinking

Abstract
Complex reasoning problems often involve implicit spatial, geometric, and structural relationships that are not explicitly encoded in text. While recent reasoning models have achieved strong performance across many domains, purely text-based reasoning struggles to represent global structural constraints in complex settings. In this paper, we introduce FigR, which integrates active visual thinking into multi-turn reasoning via end-to-end reinforcement learning. FigR externalizes intermediate structural hypotheses by constructing visual representations during problem solving. By adaptively regulating when and how visual reasoning should be invoked, FigR enables more stable and coherent reasoning over global structural properties that are difficult to capture from text alone. Experiments on challenging mathematical reasoning benchmarks demonstrate that FigR outperforms strong text-only chain-of-thought baselines. In particular, FigR improves the base model by 13.12% on AIME 2025 and 11.00% on BeyondAIME, highlighting the effectiveness of figure-guided multimodal reasoning in enhancing the stability and reliability of complex reasoning.
[Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with
Active Visual Thinking
Meiqi Chen, Fandong Meng [^2] Corresponding author., Jie Zhou WeChat AI, Tencent Inc {meiqiichen, fandongmeng, withtomzhou}@tencent.com
Figure 2: *
Figure 2: Comparison of four reasoning paradigms on a geometry problem with complex spatial constraints. Text-only CoT struggles to capture spatial relations. Unified multimodal CoT often suffers from imprecise visual grounding. Tool-augmented LVLMs are limited by predefined operations. In contrast, FigR generates executable code to construct figures, which enables geometric consistency and leads to correct reasoning.
11footnotetext: https://github.com/chenmeiqii/FIGR.
1 Introduction
Human reasoning is deeply intertwined with the ability to externalize information into visual form. When faced with complex constraints or multi-step relationships, people naturally draw diagrams, sketch intermediate states, or construct spatial layouts to reduce cognitive load and clarify underlying structure tversky2019diagrams; norman2024things. Cognitive science further shows that such diagrammatic reasoning allows people to discover patterns that are difficult or even impossible to glean from text alone donald1993origins; mandler2010spatial.
Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking illustrates how different reasoning paradigms behave when solving a geometry problem involving angle constraints, tangency conditions, and point intersections. A text-only chain-of-thought (CoT, wei2022chain) model must implicitly track a dense network of spatial relations through symbolic expressions alone. As shown in the leftmost column, this often leads to algebraic errors caused by subtle misinterpretations of positions or angles, since all geometric constraints must be internally maintained without explicit spatial grounding. To alleviate this, a natural extension is to explore unified multimodal models  team2024chameleon; wu2025janus; deng2025emerging; li2025zebra that directly generate images as part of the chain of thought. By externalizing intermediate visual representations, these models aim to make spatial relationships more explicit. However, as illustrated in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking, such directly generated images are often noisy or spatially imprecise, making it difficult to reliably encode fine-grained geometric constraints. Even minor visual inaccuracies can accumulate during multi-step reasoning and critically impair downstream symbolic computation.
In parallel, a growing line of research investigates tool-augmented large vision-language models (LVLMs), which invoke external visual tools or predefined APIs to assist reasoning, spanning prompt engineering gupta2023visual; hu2024visual; suris2023vipergpt, fine-tuning Wu_2024_CVPR; wang2025pixel, and reinforcement-learning–based approaches zheng2025deepeyes; zhang2025chain. By delegating visual processing to external tools, such methods improve controllability and enforce adherence to predefined operations. Nevertheless, as shown in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking, tool-augmented LVLMs remain restricted to a limited set of transformations applied to given images (e.g., zooming or cropping), and cannot autonomously generate task-specific visual representations from scratch. This restriction limits their expressive flexibility when reasoning requires constructing new diagrams to satisfy complex constraints.
Motivated by these limitations, we introduce FigR, which integrates visual construction directly into the reasoning loop through end-to-end reinforcement learning. Instead of generating images directly or relying on fixed toolsets, FigR produces executable code that serves as a bridge between symbolic reasoning and visual rendering. This design enables the model to actively construct and iteratively refine illustrative figures during multi-turn reasoning. The rendered figures then serve as dynamic, interpretable feedback that supports its reasoning process, much like how humans repeatedly sketch and update diagrams when solving complex problems. As illustrated in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking, this figure-guided process enforces geometric consistency and leads to correct solutions in problems where other paradigms fail. Together, this comparison highlights a key insight: effective visual reasoning requires not only access to visual representations, but also precise, controllable, and interpretable mechanisms for constructing them within the reasoning loop.
To guide this iterative reasoning–rendering process, we design an Adaptive Reward Mechanism that regulates when visual reasoning should be invoked, encouraging selective use of visual cues while discouraging unnecessary or spurious rendering behaviors. This mechanism serves as a control signal that promotes appropriate engagement with figure construction, while accuracy rewards ultimately govern reasoning quality. Crucially, this design eliminates the need for a supervised fine-tuning cold-start stage. Instead, FigR can be initialized directly from an instruction-tuned model (e.g., Qwen3-VL-32B-Instruct bai2025qwen3vl) and refined solely through reinforcement learning. Through this process, FigR autonomously learns when to invoke visual reasoning and how to integrate execution feedback into its reasoning trajectory, without relying on handcrafted heuristics or human-engineered demonstrations.
Overall, our main contributions are as follows:
  • We propose FigR, which actively conducts visual thinking during multi-turn reasoning.
  • We design a novel adaptive reward mechanism that enables selective use of visual reasoning, without requiring supervised cold-start.
  • We conduct extensive experiments across diverse mathematical reasoning benchmarks, demonstrating substantial gains over text-only baselines.
2 Related Work
2.1 Text-based Reasoning
A series of works improves upon chain‑of‑thought (CoT) prompting wei2022chain by introducing structured reasoning strategies in large language models (LLMs) achiam2023gpt; touvron2023llama. Program‑of‑Thought (PoT) chen2022program and Chain‑of‑Code li2023chain interleave natural language with executable code so that the models can offload arithmetic and symbolic manipulation to interpreters. Tree‑of‑Thought (ToT) yao2023tree and its variants represent reasoning as a branching search rather than a single chain, allowing models to explore alternative solution paths and prune unpromising branches. DeepSeek-R1 guo2025deepseek successfully applies RL to help models learn when to expand or truncate their reasoning paths. ReTool feng2025retool further teaches models when and how to call a code interpreter during mathematical problem solving, enabling dynamic interleaving of textual reasoning and tool invocation. Nevertheless, even state-of-the-art text-only models struggle with geometry, kinematics, or combinatorial tasks requiring spatial inference. This limitation stems from their inability to externalize intermediate states visually, forcing all constraints to be encoded implicitly in text—exactly the bottleneck illustrated in our introduction.
2.2 Multimodal Reasoning
Large vision–language models (LVLMs) liu2023visual; bai2025qwen2 demonstrated that LLMs can interpret images through relatively shallow alignment layers, but most of them remain limited to text-only outputs. Recent unified architectures team2024chameleon; wu2025janus; deng2025emerging; li2025zebra extend this capability by generating both text and images. However, the quality of the generated images remains limited, preventing them from supporting fine-grained and precise relational depiction. These issues become especially evident in geometric and scientific reasoning tasks, where even minor visual inaccuracies can propagate into significant logical errors.
A parallel literature explores invoking external tools, rendering engines, or predefined APIs as part of the reasoning loop. Recent work demonstrates that LVLMs can decompose a problem, apply visual operators (e.g., zooming, cropping, object detection), and interpret results to refine reasoning gupta2023visual; suris2023vipergpt; hu2024visual; shao2024visual; Wu_2024_CVPR; li2025dyfo; shen2025zoomeye; wang2025pixel. Reinforcement learning–based pipelines zheng2025deepeyes; zhang2025chain further improve tool usage efficiency by rewarding correct invocation patterns. These methods offer strong interpretability and deterministic visual transformations, making them highly effective on perception-centric benchmarks. However, they are inherently restricted to limited operations on given images. They lack support for generating new diagrams, geometric constructions, or abstract visualizations that are not directly present in the input.
3 Methodology
Figure 3: Overview of our proposed FigR. FigR alternates between textual reasoning and executable drawing steps with visual feedback. An adaptive reward mechanism dynamically regulates visual thinking based on task suitability and execution outcomes.
In this section, we describe the main components of FigR. We first outline the overall architecture and reasoning loop in Section 3.1, then describe the reinforcement-learning (RL) algorithm in Section 3.2, and finally detail the adaptive reward mechanism in Section 3.3. Figure 3 gives an overview of our FigR.
3.1 Figure-Steered Reasoning Loop
Our core idea is to embed visual construction into a model’s reasoning trajectory. Concretely, for each problem input, FigR enters a multi-turn reasoning loop in which it can interleave pure textual reasoning and executable code to generate diagrams - much like a human drawing intermediate sketches while reasoning.
Formally, at each turn $t$, FigR maintains an internal context $h_{t}$ (e.g., textual history, prior code outputs, and rendered images) and optionally a current figure/diagram $I_{t}$. The model’s policy $\pi_{\theta}$ outputs an action:
$a_{t}\sim\pi_{\theta}\left(\cdot\mid h_{t},I_{t}\right).$ (1)
Here, $a_{t}$ may be: (1) A text continuation (e.g., next reasoning step), or (2) An executable code snipper $c_{t}$ written in a general-purpose programming language (e.g., Python). If $a_{t}$ is code, the environment executes it via a sandboxed interpreter, which may produce textual outputs and render a figure:
$\big(T_{t+1},I_{t+1}\big)=\operatorname{Interpreter}\left(c_{t}\right),$ (2)
where $T_{t+1},I_{t+1}$ denotes the textual and visual feedback from code execution, respectively. The context is then updated as:
$h_{t+1}=\operatorname{UpdateContext}\left(h_{t},c_{t},T_{t+1},I_{t+1}\right).$ (3)
This multi-turn reasoning repeats until the model emits a special end token (e.g., <End>), or a maximum number of steps is reached. The produced final answer is then evaluated. Overall, the design enables both flexible visual construction and precise code-based control, providing reproducibility and interpretability.
3.2 Reinforcement Learning Procedure
To train the model to decide when to construct figures and how to integrate them into the reasoning process, we employ reinforcement learning. Specifically, we adopt the Group Relative Policy Optimization (GRPO) algorithm shao2024deepseekmath, together with a multi-turn rollout framework implemented in VeRL sheng2024hybridflow.
For each input question, GRPO samples a group of $G$ candidate trajectories (rollouts) under the current policy, computes their final rewards, and uses the group mean reward as a baseline. Trajectories that yield above-average reward are reinforced; below-average ones are suppressed. This yields sample-efficient optimization.
In our adaptation, a state includes: question, previous reasoning and code outputs, all rendered images so far, and any interpreter feedback. Action $a_{t}$ can be a text continuation or a code snippet. The environment handles code execution (if applicable), renders diagrams, and returns observations for the next turn. At the end of each rollout $\tau_{i}$ (i.e., after <End>), we compute the total reward:
$R_{i}=R_{\mathrm{acc}}\left(\tau_{i}\right)+R_{\mathrm{fmt}}\left(\tau_{i}\right)+R_{\mathrm{vis}}\left(\tau_{i}\right),$ (4)
where $R_{\text{acc }}$ measures correctness of the final answer (e.g., 1 if answer matches the ground truth, else 0), $R_{\mathrm{fmt}}$ is the format reward, granted when the model’s output adheres to the required structural format (e.g., a binary signal if the output includes separate reasoning and final answer sections), $R_{\text{vis }}$ regulates when visual reasoning is invoked (details are introduce in Section 3.3).
Within each training batch, for a given prompt, we sample a group of candidate trajectories $\left\{\tau_{i}\right\}_{i=1}^{G}$ using our policy $\pi_{\theta}$. For each trajectory, we compute $R_{i}$. We then compute the group baseline:
$\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i},$ (5)
and the relative advantage for each trajectory:
$\hat{A}_{i}=R_{i}-\bar{R}.$ (6)
GRPO updates the policy by maximizing the following surrogate objective (in expectation):
$J_{\mathrm{GRPO}}(\theta)=$ $\ \mathbb{E}_{q\sim\mathcal{D},\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|\tau_{i}|}\sum_{t=1}^{|\tau_{i}|}M_{i,t}(\theta)$ (7)
$-\beta\,D_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}),$
where
$M_{i,t}(\theta)=\min\!\left[r_{i,t}(\theta)\hat{A}_{i},\ \operatorname{clip}(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}_{i}\right].$ (8)
Here, $r_{i,t}(\theta)=\frac{\pi_{\theta}\left(a_{i,t}\mid s_{i,t}\right)}{\pi_{\theta_{\text{old }}}\left(a_{i,t}\mid s_{i,t}\right)}$ is the probability ratio; $\varepsilon$ is a clipping hyperparameter; $\pi_{\mathrm{ref}}$ is a reference policy (e.g., initial base policy) used for KL-regularization; $\beta$ is a regularization coefficient.
Above all, this objective encourages the new policy to increase the probability of trajectories whose reward is above the group-average, while penalizing divergence from the reference policy to maintain stability.
3.3 Adaptive Reward Mechanism
To encourage selective visual thinking while avoiding unnecessary drawing behaviors, we introduce an adaptive reward mechanism. Given an input question $q$, we prompt a larger language model as a classifier to predict a suitability tag $s\in\{0,1\}$, indicating whether drawing-based visual thinking is likely to be beneficial for the question.***In this work, we adopt Deepseek-V3 liu2024deepseek as the classifier because it offers a favorable balance between instruction-following capability and computational cost compared with other tested LLMs. During reinforcement learning, the visual-invocation reward is defined by jointly considering (i) whether the final answer is correct, (ii) whether the problem is suitable for visual reasoning, and (iii) whether the generated code is successfully executed. Let $R_{\mathrm{vis}}$ denote the reward associated with diagram generation, we have:
$R_{\mathrm{vis}}=\begin{cases}1.0,&\text{if }y=\hat{y},\ s=1,\ \mathrm{exec}=1,\\ 0.2,&\text{if }y=\hat{y},\ s=0,\ \mathrm{exec}=1,\\ 0,&\text{otherwise},\end{cases}$ (9)
where $y$ and $\hat{y}$ denote the ground-truth and predicted final answers, respectively, $\mathrm{exec}$ indicates whether the generated code is successfully executed and produces valid visual feedback. Under this formulation, a full visual reward (1.0) is granted only when the model answers correctly and invokes diagram generation appropriately with executable code. A reduced reward (0.2) is assigned when the answer is correct but diagram construction is deemed unnecessary, discouraging redundant visual invocations while still acknowledging successful visual thinking. All other cases, such as incorrect answers or failed executions, will receive zero reward. This design encourages selective and effective use of visual reasoning, while grounding its ultimate utility in final answer correctness.
As a result, the adaptive reward mechanism primarily serves as a control signal that guides when visual reasoning should be invoked. This design reduces unnecessary or spurious figure construction while allowing the model to learn how to integrate diagrammatic feedback effectively into reasoning.
4 Experiments
4.1 Experimental Setup
Implementation Details
For the base model, we use Qwen3-VL-32B-Instruct bai2025qwen3vl to train FigR. We set the maximum number of interaction rounds to 3, the maximum generation length to 32,768 tokens, and the sampling temperature to 0.7.
Training Dataset
We train our FigR on DeepMath-103K he2025deepmath, a rigorously decontaminated, large-scale dataset of 103,000 challenging mathematical problems with verifiable answers, explicitly curated to avoid overlap with standard test benchmarks. In addition, as described in Section 3.3, we annotate each instance in DeepMath-103K with a binary label indicating whether the current problem is suitable for figure-steered reasoning. The prompt template we use is shown in Appendix B.1.
Evaluation Datasets
To ensure comparability with prior work, we conduct evaluations on a suite of mathematical reasoning datasets that are widely used as standard benchmarks. We use AIME 2024 aime_2024, AIME 2025 aime_2025, BeyondAIME seed2025seed1, MATH 500 hendrycks2021measuring; lightman2023let, AMC amc_2023, MinervaMath lewkowycz2022solving, and the OE_TO_maths_en_COMP subset of OlympiadBench he2024olympiadbench. The dataset details are introduced in Appendix A.
Baselines
We compare FigR with the following competitive baselines:
  • Large Language Models (LLMs). These models take textual inputs and generate textual outputs. We include Qwen3-235B-A22B (Thinking) yang2025qwen3 and Qwen3-32B (Non-Thinking, Thinking) yang2025qwen3.
  • Large Vision-Language Models (LVLMs). These models could accept both text and images as inputs and produce textual outputs. We evaluate GLM-4.5V (108B) hong2025glm45v_glm41vthinking, MiMo-VL-7B-SFT-2508 yue2025mimo_vl, and Qwen3-VL-32B-Instruct.
  • Unified Multimodal Models (UMMs). These models also take both text and images as inputs, but can generate visual outputs. We include Bagel-7B-MoT deng2025emerging and Bagel-Zebra-CoT li2025zebra.
  • Tool-Augmented Vision-Language Models (TAVLMs). These models enhance visual reasoning by invoking predefined tools or APIs to process given images. We include two RL-based models DeepEyes zheng2025deepeyes and Chain-of-Focus zhang2025chain.
In addition, we provide a text-only RL baseline, where the base model is trained with GRPO solely on the problem–answer pairs from DeepMath-103K. This setting does not involve multi-turn rollout, image generation, or any external feedback, and the reward is computed only from the final textual output—covering both answer correctness and output format. For all baseline models, we set the maximum generation length, sampling temperatures, and other hyperparameters follow the recommended configurations from their respective papers or reports. The prompt templates are shown in Appendix B.2. Due to space constraints, for UMMs and TAVLMs, we report results on a subset of datasets in the main ablation study for clarity; full results are deferred to Appendix C.
Model
AIME
2024
AIME
2025
Beyond
AIME
MATH
500
AMC
Minerva
Math
Olymp
Bench
Avg.
Large Language Models
Qwen3-235B-A22B (Thinking) 83.80 80.78 52.00 95.40 93.98 46.69 74.04 75.24
Qwen3-32B (Non-Thinking) 31.00 20.20 16.00 88.60 61.45 39.34 52.08 44.10
Qwen3-32B (Thinking) 81.40 72.90 40.00 97.30 93.98 45.22 71.96 71.82
Large Vision-Language Models
GLM-4.5V (108B) 76.20 67.34 47.00 96.20 87.95 38.60 70.03 69.62
MiMo-VL-7B-SFT-2508 73.80 68.07 39.00 94.40 91.57 36.40 68.40 67.38
Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 37.87 67.51 66.55
Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 66.18 68.64
FigR (ours) 79.58 79.32 54.00 95.00 93.98 41.92 70.33 73.45
Gains +6.25 +13.12 +11.00 +1.40 +9.64 +4.05 +2.82 +6.90
Table 1: Main results (%) of FigR on seven mathematical reasoning benchmarks compared with several competitive baselines. Green-colored font indicates improvement over the baseline Qwen3-VL-32B-Instruct.
Evaluation Metrics
To ensure a stable evaluation, we default to pass@$k$ evaluation chen2021evaluating and report pass@1 metrics. Specifically, we generate 64 responses for each question of AIME 2024 and AIME 2025 (i.e., $k=64$), and $k=1$ for the remaining datasets. The pass@1 metric is then computed as: $\text{pass}@1=\frac{1}{k}\sum_{i=1}^{k}p_{i}$, where $p_{i}$ denotes the correctness of the $i$-th response.
4.2 Main Results
As shown in Table 1, FigR achieves substantial gains by incorporating visual cues through end-to-end reinforcement learning, reaching an average accuracy of 73.45% across seven mathematical benchmarks, exceeding the base model (Qwen3-VL-32B-Instruct) by 6.90% and the text-only RL baseline by 4.81%. Notably, the improvements are particularly pronounced on challenging benchmarks such as AIME 2025 and BeyondAIME. FigR significantly outperforms equivalently sized text-only models (Qwen3-32B Thinking) as well as a larger LVLM GLM-4.5V. These results indicate that purely textual chain-of-thought reasoning faces clear limitations on hard problems, whereas multimodal chain-of-thought effectively unlocks the reasoning capacity of FigR.
Figure 4: Ablation analysis across training steps. We report average response length, code count, code ratio, average code lines per executable block, and code pass rate under different settings. FigR enables more frequent and structured visual reasoning while preserving stable execution behavior, compared to ablated settings.
4.3 Ablation Study
As shown in the Table 2, we conduct ablation studies on two representative datasets: AIME 2025 and BeyondAIME, to evaluate the contributions of different components in FigR.
Prompt Engineering (PE) on the Base Model.
Prompt engineering introduces structured multimodal reasoning patterns without updating model parameters. While this leads to moderate performance gains on some datasets, the absence of a learning signal prevents the model from internalizing when and how such behaviors should be applied. As a result, the induced multimodal reasoning behaviors remain unstable, which inherently limits the attainable performance gains.
Model
AIME
2025
Beyond
AIME
Base Model 66.20 43.00
   + PE 63.33 46.00
   + SFT 53.33 38.00
   + Text-only RL 69.22 46.00
   + Bagel generated img 64.32 41.00
   + Qwen generated img 62.93 41.00
Bagel-7B-MoT 10.00 1.00
DeepEyes 2.34 0.00
FigR (ours) 79.32 52.25
   w/o ARM 70.00 49.00
   w/o img feedback 73.33 47.00
Table 2: Ablation Study on FigR. The base model is Qwen3-VL-32B-Instruct, “PE” denotes prompt engineering, “SFT” denotes supervised fine-tuning, and “ARM” denotes the adaptive reward mechanism.
Performance Comparison between SFT and RL.
When fine-tuning the base model on DeepMath-103K, supervised fine-tuning (SFT) leads to a notable performance drop (e.g., on AIME 2025, accuracy decreases from 66.20% to 53.33%). In contrast, text-only reinforcement learning yields consistent performance improvements. This indicates that SFT tends to overfit to training trajectories and generalizes poorly to unseen problems, whereas outcome-driven RL enables more robust and transferable problem-solving behavior.
Injecting Visual Information without Active Visual Reasoning.
We evaluate two baselines that directly inject visual information(+ Bagel img and + Qwen img), which generate a diagram conditioned on the problem text by using Bagel-7B-MoT and Qwen-Image wu2025qwen, respectively. Although both baselines incorporate visual inputs, neither yields consistent improvements over the base model or text-only RL. These results indicate that passively injecting images fails to provide reliable structural guidance and may instead introduce noise. This contrast highlights the importance of active, feedback-driven visual reasoning in FigR.
Performance of UMMs and TAVLMs.
Both unified multimodal models and tool-augmented vision-language models exhibit consistently weak performance. Unified models are vulnerable to cascading errors when tackling highly complex problems. Tool-augmented methods are restricted to predefined operations on given images and cannot flexibly construct task-specific diagrams. Consequently, neither paradigm provides reliable and informative visual feedback for complex mathematical reasoning. Due to space limitations, their full results are deferred to Appendix C
Ablations on FigR Components.
We further design two ablations to examine the contributions of FigR ’s components: (1) Removing the adaptive reward mechanism (ARM). We disable the visual-invocation reward, while keeping all other components unchanged. This ablation removes the explicit control signal that regulates when visual reasoning should be invoked. (2) Removing visual feedback. Building on (1), we further remove visual feedback during reinforcement learning. The model still performs multi-turn rollouts and executes code, but receives only textual feedback. This setting isolates the contribution of visual feedback in forming a stable multimodal reasoning loop. Both ablations lead to substantial performance degradation compared to FigR, demonstrating the effectiveness of the proposed components.
Figure 5: Case study of reasoning behaviors. FigR demonstrates more effective integration of visual feedback into the reasoning process, resulting in clearer intermediate reasoning and improved final answers, while the baseline models rely primarily on textual reasoning.
Evolution of active visual thinking during training.
We further analyze the emergence of visual-thinking behaviors on the BeyondAIME dataset, using four complementary metrics: response length (measured in tokens with the Qwen3-VL-32B-Instruct tokenizer), code ratio (fraction of samples that invoke code), average code lines (non-empty lines per code block), and code pass rate (fraction of code blocks that execute successfully).
(1) When visual feedback is removed (w/o vis), the model exhibits persistently low code usage despite retaining textual execution feedback. This behavior highlights a critical limitation: textual feedback fails to introduce structurally novel or irreducible information beyond what the model can already generate in natural language. Consequently, the model learns to bypass code execution and instead relies on increasingly verbose textual reasoning, as reflected by longer responses but persistently low code ratio.
(2) Removing the adaptive reward mechanism (w/o ARM) leads to a transient increase in early code usage. However, this behavior rapidly collapses: both code ratio and code length drop to zero as the model learns that indiscriminate code invocation does not improve outcomes. This pattern reflects an overuse–abandonment cycle in the absence of effective reward guidance.
(3) In contrast, FigR maintains a consistently high code ratio throughout training, accompanied by stable and substantially longer code blocks with a high execution pass rate. Visual feedback introduces an external representational modality that cannot be easily replicated within the language space, breaking the text-only reasoning loop. Consequently, code execution becomes a necessary and reliable mechanism for constructing and validating global structure, rather than an opportunistic or exploratory behavior.
4.4 Case Study
Figure 5 compares Qwen3-VL-32B-Instruct, the variant without visual feedback, and FigR on a representative example. The baseline model relies on textual reasoning and symbolic manipulation, requiring careful analytical handling of boundary cases, making it prone to subtle oversights. The variant without visual feedback introduces multi-turn rollouts and intermediate code execution, but receives only textual feedback. While this allows the model to verify intermediate numerical values or symbolic expressions, it provides no access to visual cues that reveal global structural properties of the problem (e.g., the overall shape of a function or the location of its zeros). Consequently, reasoning remains largely local and algebraic, and incorrect assumptions are difficult to detect or correct.
In contrast, FigR closes the reasoning loop by incorporating visual feedback, enabling it to observe global structural patterns and adjust subsequent reasoning steps accordingly. This leads to a more robust and interpretable solution. Overall, this case study shows that the advantage of FigR does not stem merely from executing intermediate programs, but from actively integrating visual feedback into the reasoning process. Compared with purely textual reasoning or execution with text-only feedback, active visual thinking provides complementary information that substantially improves reasoning reliability.
5 Conclusion
We introduce FigR, which integrates active visual thinking into multi-turn reasoning through end-to-end reinforcement learning. By closing the loop between reasoning and visual evidence, FigR enables models to reason over global structural properties that are difficult to infer from text alone. An adaptive reward mechanism regulates when visual reasoning should be invoked, promoting selective and reliable figure construction without unnecessary generation. Extensive experiments highlight the importance of actively incorporating visual feedback during reasoning and suggest a promising direction for interpretable multimodal reasoning.
Model
AIME
2024
AIME
2025
Beyond
AIME
MATH
500
AMC
Minerva
Math
Olymp
Bench
Avg.
Large Language Models
Qwen3-235B-A22B (Thinking) 83.80 80.78 52.00 95.40 93.98 46.69 74.04 75.24
Qwen3-32B (Non-Thinking) 31.00 20.20 16.00 88.60 61.45 39.34 52.08 44.10
Qwen3-32B (Thinking) 81.40 72.90 40.00 97.30 93.98 45.22 71.96 71.82
Large Vision-Language Models
GLM-4.5V (108B) 76.20 67.34 47.00 96.20 87.95 38.60 70.03 69.62
MiMo-VL-7B-SFT-2508 73.80 68.07 39.00 94.40 91.57 36.40 68.40 67.38
Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 37.87 67.51 66.55
Unified Multimodal Models
Bagel-7B-MoT 10.00 3.33 1.00 56.00 22.89 9.56 26.11 18.70
Bagel-Zebra-CoT (7B) 0.00 5.55 0.00 20.60 9.64 2.94 8.46 6.74
Tool-Augmented Vision-Language Models
DeepEyes 6.98 2.34 0.00 11.00 18.07 5.15 13.65 8.17
Chain-of-Focus 1.30 0.00 0.00 1.20 0.00 0.74 0.45 0.53
Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 66.18 68.64
FigR (ours) 79.58 79.32 54.00 95.00 93.98 41.92 70.33 73.45
Gains +6.25 +13.12 +11.00 +1.40 +9.64 +4.05 +2.82 +6.90
Table 3: Full experimental results (%) on seven mathematical reasoning benchmarks. Green-colored font indicates the improvement of FigR over the baseline Qwen3-VL-32B-Instruct.
Figure 6: Prompt template for the figure-steered suitability classifier.
Appendix A Dataset Details
We evaluate models on a suite of challenging mathematical reasoning benchmarks, spanning competition problems, standardized mathematics datasets, and advanced reasoning tasks:
(1) AIME 2024 aime_2024 dataset contains 30 problems from the American Invitational Mathematics Examination 2024. Each problem requires multi-step reasoning across algebra, number theory, combinatorics, geometry, and probability.
(2) AIME 2025 aime_2025. Like the AIME 2024 dataset, the AIME 2025 dataset comprises the 30 problems from the 2025 AIME competitions.
(3) Beyond AIME seed2025seed1 is a recently proposed benchmark of 100 problems. It is designed to push beyond standard AIME problems by emphasizing questions that reduce memorization and increase the complexity of reasoning.
(4) MATH 500 hendrycks2021measuring; lightman2023let. The MATH dataset contains 12,500 problems with step-by-step solutions spanning diverse mathematical topics. For evaluation focused on high-difficulty items, the MATH 500 subset (500 problems) is often used.
(5) AMC amc_2023. The American Mathematics Competitions (AMC) dataset consists of 83 problems from the AMC series, a set of standardized mathematics contests administered annually.
(6) MinervaMath lewkowycz2022solving includes 272 undergraduate-level quantitative reasoning problems that require logical deductions and multi-step solutions, serving as a more advanced testbed beyond high-school competition problems.
(7) OlympiadBench he2024olympiadbench is an Olympiad-level bilingual multimodal scientific benchmark that includes thousands of challenging mathematics and physics problems compiled from international competitions, with detailed annotations for step-by-step reasoning. In this work, we specifically evaluate the math-focused English comprehensive subsets from OlympiadBench: OE_TO_maths_en_COMP. This subset is an English open-ended, text-only, comprehensive math subset containing 674 problems, emphasizing rigorous mathematical reasoning in text form.
Appendix B Prompt Templates
B.1 Prompt Template for Suitability Classifier
We provide a dedicated prompt template for the figure-steered suitability classifier as introduced in Section 3.3 and Section 4.1, which is used to determine whether a given problem is suitable for visual-guided reasoning. As shown in Figure 6, the template instructs the model to assess the necessity and usefulness of visual sketches based solely on the problem statement.
B.2 Prompt Template for Reasoning
We adopt three different categories of prompt templates corresponding to the reasoning settings. (1) As shown in Figure 7, a multi-turn prompt with drawing feedback is used for FigR, the variant without visual feedback, and the variant without ARM. Models can iteratively generate output and receive feedback from the interpreter. (2) As shown in Figure 8, a standard single-turn prompt is adopted for text-only RL and other baselines, requiring the model to directly output the final answer without intermediate interactions. (3) For DeepEyes and Chain-of-Focus, we adopt the prompt templates recommended by the original papers.
Figure 7: Prompt template for multi-turn CoT.
Figure 8: Prompt template for single-turn CoT.
Appendix C Additional Experimental Results
As shown in Table 3, we present the full experimental results, including unified multimodal models and tool-augmented vision-language models that are partially reported in the main context.

### v2

@@
-Active Visual Thinking
+Executable Visual States
 Abstract
-Complex reasoning problems often involve implicit spatial, geometric, and structural relationships that are not explicitly encoded in text.
-While recent reasoning models have achieved strong performance across many domains, purely text-based reasoning struggles to represent global structural constraints in complex settings.
-In this paper, we introduce FigR, which integrates active visual thinking into multi-turn reasoning via end-to-end reinforcement learning.
-FigR externalizes intermediate structural hypotheses by constructing visual representations during problem solving.
-By adaptively regulating when and how visual reasoning should be invoked, FigR enables more stable and coherent reasoning over global structural properties that are difficult to capture from text alone.
-Experiments on challenging mathematical reasoning benchmarks demonstrate that FigR outperforms strong text-only chain-of-thought baselines.
-In particular, FigR improves the base model by 13.12% on AIME 2025 and 11.00% on BeyondAIME, highlighting the effectiveness of figure-guided multimodal reasoning in enhancing the stability and reliability of complex reasoning.
+Complex reasoning problems often involve implicit spatial and geometric relationships that are not explicitly encoded in text.
+While recent reasoning models perform well across many domains, purely text-based reasoning struggles to capture structural constraints in complex settings.
+In this paper, we introduce FigR, which integrates executable visual construction into multi-turn reasoning via end-to-end reinforcement learning.
+Rather than relying solely on textual chains of thought, FigR externalizes intermediate hypotheses by generating executable code that constructs diagrams within the reasoning loop.
+An adaptive reward mechanism selectively regulates when visual construction is invoked, enabling more consistent reasoning over latent global properties that are difficult to infer from text alone.
+Experiments on eight challenging mathematical benchmarks demonstrate that FigR outperforms strong text-only chain-of-thought baselines, improving the base model by 13.12% on AIME 2025 and 11.00% on BeyondAIME.
+These results highlight the effectiveness of precise, controllable figure construction of FigR in enhancing complex reasoning ability.
 [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with
-Active Visual Thinking
+Executable Visual States
 Meiqi Chen, Fandong Meng [^2] Corresponding author., Jie Zhou WeChat AI, Tencent Inc {meiqiichen, fandongmeng, withtomzhou}@tencent.com
 Figure 2: *
-Figure 2: Comparison of four reasoning paradigms on a geometry problem with complex spatial constraints.
-Text-only CoT struggles to capture spatial relations.
-Unified multimodal CoT often suffers from imprecise visual grounding.
-Tool-augmented LVLMs are limited by predefined operations.
-In contrast, FigR generates executable code to construct figures, which enables geometric consistency and leads to correct reasoning.
+Figure 2: Comparison of reasoning paradigms by intermediate state representation.
+Text-only reasoning relies on implicit symbolic states, unified multimodal models generate uncontrolled visual states, and tool-augmented models operate on given images with predefined transformations.
+In contrast, FigR constructs executable visual states within the reasoning loop, enabling precise, revisable diagrams that enforce geometric consistency.
 11footnotetext: https://github.com/chenmeiqii/FIGR.
@@
 When faced with complex constraints or multi-step relationships, people naturally draw diagrams, sketch intermediate states, or construct spatial layouts to reduce cognitive load and clarify underlying structure tversky2019diagrams; norman2024things.
-Cognitive science further shows that such diagrammatic reasoning allows people to discover patterns that are difficult or even impossible to glean from text alone donald1993origins; mandler2010spatial.
-Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking illustrates how different reasoning paradigms behave when solving a geometry problem involving angle constraints, tangency conditions, and point intersections.
-A text-only chain-of-thought (CoT, wei2022chain) model must implicitly track a dense network of spatial relations through symbolic expressions alone.
-As shown in the leftmost column, this often leads to algebraic errors caused by subtle misinterpretations of positions or angles, since all geometric constraints must be internally maintained without explicit spatial grounding.
-To alleviate this, a natural extension is to explore unified multimodal models  team2024chameleon; wu2025janus; deng2025emerging; li2025zebra that directly generate images as part of the chain of thought.
-By externalizing intermediate visual representations, these models aim to make spatial relationships more explicit.
-However, as illustrated in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking, such directly generated images are often noisy or spatially imprecise, making it difficult to reliably encode fine-grained geometric constraints.
-Even minor visual inaccuracies can accumulate during multi-step reasoning and critically impair downstream symbolic computation.
-In parallel, a growing line of research investigates tool-augmented large vision-language models (LVLMs), which invoke external visual tools or predefined APIs to assist reasoning, spanning prompt engineering gupta2023visual; hu2024visual; suris2023vipergpt, fine-tuning Wu_2024_CVPR; wang2025pixel, and reinforcement-learning–based approaches zheng2025deepeyes; zhang2025chain.
-By delegating visual processing to external tools, such methods improve controllability and enforce adherence to predefined operations.
-Nevertheless, as shown in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking, tool-augmented LVLMs remain restricted to a limited set of transformations applied to given images (e.g., zooming or cropping), and cannot autonomously generate task-specific visual representations from scratch.
-This restriction limits their expressive flexibility when reasoning requires constructing new diagrams to satisfy complex constraints.
-Motivated by these limitations, we introduce FigR, which integrates visual construction directly into the reasoning loop through end-to-end reinforcement learning.
-Instead of generating images directly or relying on fixed toolsets, FigR produces executable code that serves as a bridge between symbolic reasoning and visual rendering.
-This design enables the model to actively construct and iteratively refine illustrative figures during multi-turn reasoning.
-The rendered figures then serve as dynamic, interpretable feedback that supports its reasoning process, much like how humans repeatedly sketch and update diagrams when solving complex problems.
-As illustrated in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Active Visual Thinking, this figure-guided process enforces geometric consistency and leads to correct solutions in problems where other paradigms fail.
-Together, this comparison highlights a key insight: effective visual reasoning requires not only access to visual representations, but also precise, controllable, and interpretable mechanisms for constructing them within the reasoning loop.
-To guide this iterative reasoning–rendering process, we design an Adaptive Reward Mechanism that regulates when visual reasoning should be invoked, encouraging selective use of visual cues while discouraging unnecessary or spurious rendering behaviors.
-This mechanism serves as a control signal that promotes appropriate engagement with figure construction, while accuracy rewards ultimately govern reasoning quality.
-Crucially, this design eliminates the need for a supervised fine-tuning cold-start stage.
-Instead, FigR can be initialized directly from an instruction-tuned model (e.g., Qwen3-VL-32B-Instruct bai2025qwen3vl) and refined solely through reinforcement learning.
-Through this process, FigR autonomously learns when to invoke visual reasoning and how to integrate execution feedback into its reasoning trajectory, without relying on handcrafted heuristics or human-engineered demonstrations.
+Cognitive science further suggests that such diagrammatic reasoning allows people to discover patterns that are difficult or even impossible to glean from text alone donald1993origins; mandler2010spatial.
+Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Executable Visual States provides an illustrative comparison of how different reasoning paradigms behave on a geometry problem involving angle constraints, tangency conditions, and point intersections.
+In text-only chain-of-thought (CoT, wei2022chain) models, all spatial relations must be implicitly maintained through symbolic expressions.
+This places a heavy burden on internal representations and often leads to cascading algebraic errors when subtle geometric constraints are misinterpreted.
+A natural extension is to consider unified multimodal models  team2024chameleon; wu2025janus; deng2025emerging; li2025zebra, which generate images as part of the reasoning process to make spatial relationships explicit.
+While conceptually appealing, these models lack precise control over the generated visual content.
+Since image generation is not grounded in executable constraints, even small spatial inconsistencies can propagate across reasoning steps, limiting their reliability in problems that require fine-grained geometric precision.
+In parallel, tool-augmented large vision-language models (LVLMs) incorporate external visual tools or predefined APIs to assist reasoning, spanning prompt engineering gupta2023visual; hu2024visual; suris2023vipergpt, fine-tuning Wu_2024_CVPR; wang2025pixel, or reinforcement-learning zheng2025deepeyes; zhang2025chain.
+By delegating visual operations to explicit tools, these approaches improve controllability and execution fidelity.
+However, they are fundamentally constrained to operating on given images and predefined transformations (e.g., zooming or cropping), which prevents them from autonomously constructing task-specific diagrams required by complex reasoning problems.
+Motivated by these limitations, we introduce FigR, which integrates visual construction directly into the reasoning loop via end-to-end reinforcement learning.
+Rather than relying on unconstrained image generation or fixed toolsets, FigR produces executable code that bridges symbolic reasoning and visual rendering.
+This allows the model to actively construct and iteratively refine diagrams during multi-turn inference, using rendered figures as explicit, stateful feedback—analogous to how humans repeatedly sketch and update diagrams when solving complex problems.
+As illustrated in Figure [Uncaptioned image] Figure It Out: Improve the Frontier of Reasoning with Executable Visual States, this mechanism enforces geometric consistency and supports correct reasoning in cases where other paradigms fail.
+To regulate this reasoning–rendering process, we design an Adaptive Reward Mechanism that selectively encourages visual construction when it is beneficial, while discouraging unnecessary or spurious rendering behaviors.
+Importantly, this design removes the need for a supervised visual reasoning cold-start stage.
+Instead, FigR can be initialized from an instruction-tuned model (e.g., Qwen3-VL-32B-Instruct bai2025qwen3vl) and refined purely through reinforcement learning.
+Through this process, FigR autonomously learns both when to invoke visual reasoning and how to integrate execution feedback into its reasoning trajectory, without relying on task-specific supervised visual reasoning data.
 Overall, our main contributions are as follows:
   • We propose FigR, which actively conducts visual thinking during multi-turn reasoning.
-  • We design a novel adaptive reward mechanism that enables selective use of visual reasoning, without requiring supervised cold-start.
-  • We conduct extensive experiments across diverse mathematical reasoning benchmarks, demonstrating substantial gains over text-only baselines.
+  • We introduce an adaptive reward mechanism that regulates the selective use of visual reasoning without requiring supervised cold-start.
+  • We demonstrate substantial improvements over strong text-only baselines across diverse mathematical reasoning benchmarks.
 2 Related Work
-2.1 Text-based Reasoning
-A series of works improves upon chain‑of‑thought (CoT) prompting wei2022chain by introducing structured reasoning strategies in large language models (LLMs) achiam2023gpt; touvron2023llama.
-Program‑of‑Thought (PoT) chen2022program and Chain‑of‑Code li2023chain interleave natural language with executable code so that the models can offload arithmetic and symbolic manipulation to interpreters.
-Tree‑of‑Thought (ToT) yao2023tree and its variants represent reasoning as a branching search rather than a single chain, allowing models to explore alternative solution paths and prune unpromising branches.
-DeepSeek-R1 guo2025deepseek successfully applies RL to help models learn when to expand or truncate their reasoning paths.
-ReTool feng2025retool further teaches models when and how to call a code interpreter during mathematical problem solving, enabling dynamic interleaving of textual reasoning and tool invocation.
-Nevertheless, even state-of-the-art text-only models struggle with geometry, kinematics, or combinatorial tasks requiring spatial inference.
-This limitation stems from their inability to externalize intermediate states visually, forcing all constraints to be encoded implicitly in text—exactly the bottleneck illustrated in our introduction.
-2.2 Multimodal Reasoning
-Large vision–language models (LVLMs) liu2023visual; bai2025qwen2 demonstrated that LLMs can interpret images through relatively shallow alignment layers, but most of them remain limited to text-only outputs.
-Recent unified architectures team2024chameleon; wu2025janus; deng2025emerging; li2025zebra extend this capability by generating both text and images.
-However, the quality of the generated images remains limited, preventing them from supporting fine-grained and precise relational depiction.
-These issues become especially evident in geometric and scientific reasoning tasks, where even minor visual inaccuracies can propagate into significant logical errors.
-A parallel literature explores invoking external tools, rendering engines, or predefined APIs as part of the reasoning loop.
-Recent work demonstrates that LVLMs can decompose a problem, apply visual operators (e.g., zooming, cropping, object detection), and interpret results to refine reasoning gupta2023visual; suris2023vipergpt; hu2024visual; shao2024visual; Wu_2024_CVPR; li2025dyfo; shen2025zoomeye; wang2025pixel.
-Reinforcement learning–based pipelines zheng2025deepeyes; zhang2025chain further improve tool usage efficiency by rewarding correct invocation patterns.
-These methods offer strong interpretability and deterministic visual transformations, making them highly effective on perception-centric benchmarks.
-However, they are inherently restricted to limited operations on given images.
-They lack support for generating new diagrams, geometric constructions, or abstract visualizations that are not directly present in the input.
+2.1 Text-based and Programmatic Reasoning
+A series of works extends chain-of-thought (CoT) prompting wei2022chain by introducing structured reasoning strategies for large language models (LLMs) achiam2023gpt; touvron2023llama.
+Program-of-Thought (PoT) chen2022program and Chain-of-Code li2023chain interleave natural language with executable code, allowing models to offload arithmetic and symbolic manipulation to external interpreters.
+Tree-of-Thought (ToT) yao2023tree and its variants further represent reasoning as a branching search process, enabling exploration of alternative solution paths and pruning of unpromising branches.
+More recently, reinforcement learning has been applied to regulate reasoning behaviors, such as learning when to expand or truncate reasoning trajectories guo2025deepseek, or when and how to invoke a code interpreter during problem solving feng2025retool.
+Despite these advances, such approaches operate primarily over textual or symbolic representations.
+Intermediate reasoning states remain implicit and must be maintained internally by the model, which poses fundamental challenges for tasks involving geometry, kinematics, or other forms of spatial reasoning.
+Without explicit external representations, complex relational constraints are prone to the accumulation of errors, limiting the reliability of purely text-based and programmatic reasoning in structurally demanding settings.
+2.2 Unconstrained Visual Generation for Reasoning
+Large vision-language models (LVLMs) liu2023visual; bai2025qwen2 demonstrate that LLMs can interpret images through relatively shallow alignment layers, but most of them remain limited to text-only outputs.
+Recent unified multimodal architectures team2024chameleon; wu2025janus; deng2025emerging; li2025zebra further extend generation to both text and images within a single model, enabling visual content to be produced as part of the reasoning process.
+While conceptually appealing, these approaches rely on unconstrained image generation without explicit mechanisms to enforce geometric precision or relational consistency.
+As a result, the generated visual representations are often noisy or spatially imprecise, limiting their effectiveness as reliable intermediate states for reasoning.
+This limitation becomes particularly pronounced in mathematical and scientific domains, where even minor visual inaccuracies can propagate across reasoning steps and lead to significant downstream errors.
+2.3 Tool-Augmented Visual Reasoning
+Another line of research augments LVLMs with external visual tools, rendering engines, or predefined APIs.
+These methods decompose a problem into sub-steps, apply visual operators such as zooming, cropping, or drawing auxiliary lines, and interpret the results to guide subsequent reasoning gupta2023visual; suris2023vipergpt; hu2024visual; shao2024visual; Wu_2024_CVPR; li2025dyfo; shen2025zoomeye; wang2025pixel.
+Reinforcement learning-based pipelines further improve tool usage efficiency by explicitly rewarding appropriate invocation patterns zheng2025deepeyes; zhang2025chain.
+Tool-augmented visual reasoning approaches provide strong controllability, as their visual operations are deterministic and explicitly defined.
+However, they are inherently restricted to operating on given images and a fixed set of predefined transformations.
+Consequently, they lack the ability to autonomously construct new diagrams, geometric configurations, or abstract visual representations that are not present in the input.
+This restriction limits their applicability in reasoning tasks that require dynamic generation and refinement of task-specific visual structures.
 3 Methodology
-Figure 3: Overview of our proposed FigR.
-FigR alternates between textual reasoning and executable drawing steps with visual feedback.
-An adaptive reward mechanism dynamically regulates visual thinking based on task suitability and execution outcomes.
+Figure 3: Overview of FigR.
+FigR alternates between textual reasoning and executable visual construction within a unified reasoning loop.
+An adaptive reward mechanism selectively regulates when visual construction is invoked, based on question suitability, execution outcomes, and final answer correctness.
 In this section, we describe the main components of FigR.
-We first outline the overall architecture and reasoning loop in Section 3.1, then describe the reinforcement-learning (RL) algorithm in Section 3.2, and finally detail the adaptive reward mechanism in Section 3.3.
-Figure 3 gives an overview of our FigR.
-3.1 Figure-Steered Reasoning Loop
-Our core idea is to embed visual construction into a model’s reasoning trajectory.
-Concretely, for each problem input, FigR enters a multi-turn reasoning loop in which it can interleave pure textual reasoning and executable code to generate diagrams - much like a human drawing intermediate sketches while reasoning.
-Formally, at each turn $t$, FigR maintains an internal context $h_{t}$ (e.g., textual history, prior code outputs, and rendered images) and optionally a current figure/diagram $I_{t}$.
-The model’s policy $\pi_{\theta}$ outputs an action:
-$a_{t}\sim\pi_{\theta}\left(\cdot\mid h_{t},I_{t}\right).$ (1)
-Here, $a_{t}$ may be: (1) A text continuation (e.g., next reasoning step), or (2) An executable code snipper $c_{t}$ written in a general-purpose programming language (e.g., Python).
-If $a_{t}$ is code, the environment executes it via a sandboxed interpreter, which may produce textual outputs and render a figure:
+Section 3.1 introduces the executable reasoning loop, Section 3.2 details the reinforcement learning procedure, and Section 3.3 presents the adaptive reward mechanism.
+3.1 Executable Visual Construction within the Reasoning Loop
+FigR integrates executable visual construction directly into its multi-turn reasoning process.
+For each input, the model interleaves textual inference and code generation to construct diagrams, analogous to how humans iteratively sketch intermediate states while reasoning.
+Formally, at each reasoning step $t$, the model maintains a state consisting of a textual context $h_{t}$ (including prior reasoning steps, code outputs, and interpreter feedback) and an optional rendered diagram $I_{t}$.
+The policy $\pi_{\theta}$ samples an action:
+$a_{t}\sim\pi_{\theta}\left(\cdot\mid h_{t},I_{t}\right),$ (1)
+where $a_{t}$ is either (i) a textual continuation or (ii) a code snippet $c_{t}$ generated by the policy itself.
+When a code action is emitted, it is executed by a sandboxed interpreter, which may produce textual feedback $T_{t+1}$ and/or a rendered diagram $I_{t+1}$:
 $\big(T_{t+1},I_{t+1}\big)=\operatorname{Interpreter}\left(c_{t}\right),$ (2)
-where $T_{t+1},I_{t+1}$ denotes the textual and visual feedback from code execution, respectively. The context is then updated as:
+The context is then updated accordingly:
 $h_{t+1}=\operatorname{UpdateContext}\left(h_{t},c_{t},T_{t+1},I_{t+1}\right).$ (3)
-This multi-turn reasoning repeats until the model emits a special end token (e.g., <End>), or a maximum number of steps is reached.
-The produced final answer is then evaluated.
-Overall, the design enables both flexible visual construction and precise code-based control, providing reproducibility and interpretability.
-3.2 Reinforcement Learning Procedure
-To train the model to decide when to construct figures and how to integrate them into the reasoning process, we employ reinforcement learning.
-Specifically, we adopt the Group Relative Policy Optimization (GRPO) algorithm shao2024deepseekmath, together with a multi-turn rollout framework implemented in VeRL sheng2024hybridflow.
-For each input question, GRPO samples a group of $G$ candidate trajectories (rollouts) under the current policy, computes their final rewards, and uses the group mean reward as a baseline.
-Trajectories that yield above-average reward are reinforced; below-average ones are suppressed.
-This yields sample-efficient optimization.
-In our adaptation, a state includes: question, previous reasoning and code outputs, all rendered images so far, and any interpreter feedback.
-Action $a_{t}$ can be a text continuation or a code snippet.
-The environment handles code execution (if applicable), renders diagrams, and returns observations for the next turn.
-At the end of each rollout $\tau_{i}$ (i.e., after <End>), we compute the total reward:
-$R_{i}=R_{\mathrm{acc}}\left(\tau_{i}\right)+R_{\mathrm{fmt}}\left(\tau_{i}\right)+R_{\mathrm{vis}}\left(\tau_{i}\right),$ (4)
-where $R_{\text{acc }}$ measures correctness of the final answer (e.g., 1 if answer matches the ground truth, else 0), $R_{\mathrm{fmt}}$ is the format reward, granted when the model’s output adheres to the required structural format (e.g., a binary signal if the output includes separate reasoning and final answer sections), $R_{\text{vis }}$ regulates when visual reasoning is invoked (details are introduce in Section 3.3).
-Within each training batch, for a given prompt, we sample a group of candidate trajectories $\left\{\tau_{i}\right\}_{i=1}^{G}$ using our policy $\pi_{\theta}$.
-For each trajectory, we compute $R_{i}$.
-We then compute the group baseline:
-$\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i},$ (5)
-and the relative advantage for each trajectory:
-$\hat{A}_{i}=R_{i}-\bar{R}.$ (6)
-GRPO updates the policy by maximizing the following surrogate objective (in expectation):
+This process repeats until a termination token is emitted or a step limit is reached.
+By representing diagrams as executable and reproducible states, FigR enables precise and controllable visual feedback throughout the reasoning process.
+3.2 Reinforcement Learning with Multi-turn Executable Reasoning
+Training FigR requires learning long-horizon reasoning policies that interleave textual inference with executable visual construction.
+In this setting, the quality of individual actions (e.g., whether to emit code at a given step) cannot be evaluated locally, as their utility depends on the final reasoning outcome and the resulting visual states.
+We therefore formulate learning at the trajectory level and employ reinforcement learning to optimize multi-turn reasoning behaviors.
+We adopt Group Relative Policy Optimization (GRPO) shao2024deepseekmath, which is well-suited for trajectory-level supervision with delayed terminal rewards.
+For each input question, GRPO samples a group of $G$ candidate trajectories under the current policy.
+Each trajectory $\tau_{i}$ consists of a sequence of textual and code actions, together with their executed interpreter feedback, and terminates when the model emits a special end token.
+After rollout termination, we compute a scalar reward:
+$R_{i}=R_{\mathrm{acc}}(\tau_{i})+R_{\mathrm{fmt}}(\tau_{i})+R_{\mathrm{vis}}(\tau_{i}),$ (4)
+which jointly reflects final answer correctness, output format compliance, and the appropriateness of visual construction (Section 3.3).
+Importantly, rewards are assigned only at the trajectory level, as intermediate actions contribute to reasoning quality through their cumulative effects.
+To stabilize learning under such delayed supervision, GRPO compares trajectories generated from the same prompt and uses the group-average reward as a baseline:
+$\bar{R}=\frac{1}{G}\sum_{i=1}^{G}R_{i}.$ (5)
+The relative advantage of each trajectory is then defined as:
+$\hat{A}_{i}=R_{i}-\bar{R},$ (6)
+which provides a low-variance learning signal that emphasizes relative behavioral differences, such as when executable visual construction leads to better outcomes than purely textual reasoning.
+The policy is updated by maximizing the following surrogate objective:
 $J_{\mathrm{GRPO}}(\theta)=$ $\ \mathbb{E}_{q\sim\mathcal{D},\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|\tau_{i}|}\sum_{t=1}^{|\tau_{i}|}M_{i,t}(\theta)$ (7)
@@
 $M_{i,t}(\theta)=\min\!\left[r_{i,t}(\theta)\hat{A}_{i},\ \operatorname{clip}(r_{i,t}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}_{i}\right].$ (8)
-Here, $r_{i,t}(\theta)=\frac{\pi_{\theta}\left(a_{i,t}\mid s_{i,t}\right)}{\pi_{\theta_{\text{old }}}\left(a_{i,t}\mid s_{i,t}\right)}$ is the probability ratio; $\varepsilon$ is a clipping hyperparameter; $\pi_{\mathrm{ref}}$ is a reference policy (e.g., initial base policy) used for KL-regularization; $\beta$ is a regularization coefficient.
-Above all, this objective encourages the new policy to increase the probability of trajectories whose reward is above the group-average, while penalizing divergence from the reference policy to maintain stability.
+Here, $r_{i,t}(\theta)=\frac{\pi_{\theta}(a_{i,t}\mid s_{i,t})}{\pi_{\theta_{\text{old}}}(a_{i,t}\mid s_{i,t})}$ denotes the probability ratio between the updated and previous policies, $\varepsilon$ is the clipping threshold, and $\pi_{\mathrm{ref}}$ is a reference policy used for KL regularization.
+This objective encourages the policy to increase the likelihood of trajectories that achieve higher relative rewards, while maintaining stability during optimization.
+Overall, GRPO enables FigR to learn multi-turn reasoning strategies in which the decision to invoke visual construction and the resulting reasoning quality are optimized jointly, without requiring step-wise supervision or an explicit value model.
 3.3 Adaptive Reward Mechanism
-To encourage selective visual thinking while avoiding unnecessary drawing behaviors, we introduce an adaptive reward mechanism.
-Given an input question $q$, we prompt a larger language model as a classifier to predict a suitability tag $s\in\{0,1\}$, indicating whether drawing-based visual thinking is likely to be beneficial for the question.***In this work, we adopt Deepseek-V3 liu2024deepseek as the classifier because it offers a favorable balance between instruction-following capability and computational cost compared with other tested LLMs.
-During reinforcement learning, the visual-invocation reward is defined by jointly considering (i) whether the final answer is correct, (ii) whether the problem is suitable for visual reasoning, and (iii) whether the generated code is successfully executed.
-Let $R_{\mathrm{vis}}$ denote the reward associated with diagram generation, we have:
+While executable visual construction can significantly benefit certain reasoning problems, indiscriminate invocation of diagram generation may introduce unnecessary complexity or distract from effective textual reasoning.
+To address this trade-off, we introduce an adaptive reward mechanism that acts as a policy-level control signal, guiding when visual construction should be invoked during multi-turn reasoning.
+For each input question, we first estimate whether diagrammatic reasoning is likely to be beneficial by querying an auxiliary language model classifier, which outputs a binary suitability label $s\in\{0,1\}$.
+Importantly, this classifier does not provide step-wise supervision; instead, it modulates reward magnitude, encouraging selective use of visual construction and making the policy robust to moderate misclassification.
+During reinforcement learning, the visual-invocation reward $R_{\mathrm{vis}}$ is evaluated only at the trajectory level and is coupled with final answer correctness.
+Specifically, a positive reward is granted when executable visual construction contributes to a correct solution, with higher weight assigned when diagrammatic reasoning is predicted to be appropriate for the task:
 $R_{\mathrm{vis}}=\begin{cases}1.0,&\text{if }y=\hat{y},\ s=1,\ \mathrm{exec}=1,\\ 0.2,&\text{if }y=\hat{y},\ s=0,\ \mathrm{exec}=1,\\ 0,&\text{otherwise},\end{cases}$ (9)
-where $y$ and $\hat{y}$ denote the ground-truth and predicted final answers, respectively, $\mathrm{exec}$ indicates whether the generated code is successfully executed and produces valid visual feedback.
-Under this formulation, a full visual reward (1.0) is granted only when the model answers correctly and invokes diagram generation appropriately with executable code.
-A reduced reward (0.2) is assigned when the answer is correct but diagram construction is deemed unnecessary, discouraging redundant visual invocations while still acknowledging successful visual thinking.
-All other cases, such as incorrect answers or failed executions, will receive zero reward.
-This design encourages selective and effective use of visual reasoning, while grounding its ultimate utility in final answer correctness.
-As a result, the adaptive reward mechanism primarily serves as a control signal that guides when visual reasoning should be invoked.
-This design reduces unnecessary or spurious figure construction while allowing the model to learn how to integrate diagrammatic feedback effectively into reasoning.
+Here, $y$ and $\hat{y}$ denote the ground-truth and predicted final answers, respectively, and $\mathrm{exec}$ indicates successful execution of the generated code.
+Rather than explicitly supervising individual drawing actions, this reward formulation biases the policy toward reasoning strategies in which visual construction is used purposefully and effectively.
+By tying visual rewards to overall task success, the model learns to treat diagram construction as a meaningful intermediate state rather than a mandatory or habitual operation.
+This design discourages redundant or spurious visual actions while preserving the flexibility to exploit executable visual feedback when it improves reasoning outcomes.
 4 Experiments
@@
 Implementation Details
-For the base model, we use Qwen3-VL-32B-Instruct bai2025qwen3vl to train FigR.
+We conduct reinforcement learning using the VeRL framework sheng2024hybridflow, with Qwen3-VL-32B-Instruct bai2025qwen3vl as the base policy.
 We set the maximum number of interaction rounds to 3, the maximum generation length to 32,768 tokens, and the sampling temperature to 0.7.
 Training Dataset
-We train our FigR on DeepMath-103K he2025deepmath, a rigorously decontaminated, large-scale dataset of 103,000 challenging mathematical problems with verifiable answers, explicitly curated to avoid overlap with standard test benchmarks.
-In addition, as described in Section 3.3, we annotate each instance in DeepMath-103K with a binary label indicating whether the current problem is suitable for figure-steered reasoning.
-The prompt template we use is shown in Appendix B.1.
+We train FigR on DeepMath-103K he2025deepmath, a rigorously decontaminated dataset of 103,000 challenging mathematical problems with verifiable answers.
+Following Section 3.3, we annotate each instance with a binary suitability label indicating whether executable visual construction is likely to be beneficial.
+An auxiliary language model (Deepseek-V3 liu2024deepseek) is used for this annotation; prompt details are provided in Appendix B.1.
 Evaluation Datasets
@@
 We compare FigR with the following competitive baselines:
-  • Large Language Models (LLMs).
-These models take textual inputs and generate textual outputs.
-We include Qwen3-235B-A22B (Thinking) yang2025qwen3 and Qwen3-32B (Non-Thinking, Thinking) yang2025qwen3.
-  • Large Vision-Language Models (LVLMs).
-These models could accept both text and images as inputs and produce textual outputs.
-We evaluate GLM-4.5V (108B) hong2025glm45v_glm41vthinking, MiMo-VL-7B-SFT-2508 yue2025mimo_vl, and Qwen3-VL-32B-Instruct.
-  • Unified Multimodal Models (UMMs). These models also take both text and images as inputs, but can generate visual outputs. We include Bagel-7B-MoT deng2025emerging and Bagel-Zebra-CoT li2025zebra.
+  • Implicit-State Reasoning Models .
+These models generate text-only outputs that implicitly maintain all intermediate states.
+We include three LLMs: Qwen3-235B-A22B (Thinking) yang2025qwen3 and Qwen3-32B (Non-Thinking, Thinking) yang2025qwen3; and three LVLMs: GLM-4.5V (108B), Qwen3-VL-8B-Instruct, and Qwen3-VL-32B-Instruct.
+  • Unconstrained Visual Generation Models.
+Unified multimodal models (UMMs) that generate images as part of the reasoning process, including Bagel-7B-MoT deng2025emerging and Bagel-Zebra-CoT li2025zebra.
   • Tool-Augmented Vision-Language Models (TAVLMs).
-These models enhance visual reasoning by invoking predefined tools or APIs to process given images.
-We include two RL-based models DeepEyes zheng2025deepeyes and Chain-of-Focus zhang2025chain.
-In addition, we provide a text-only RL baseline, where the base model is trained with GRPO solely on the problem–answer pairs from DeepMath-103K.
-This setting does not involve multi-turn rollout, image generation, or any external feedback, and the reward is computed only from the final textual output—covering both answer correctness and output format.
+Models that rely on predefined visual tools or APIs operating on given images, including DeepEyes zheng2025deepeyes and Chain-of-Focus zhang2025chain.
+We additionally include a text-only RL baseline trained with GRPO on DeepMath-103K, which optimizes trajectory-level correctness and format rewards without executable visual construction or interpreter feedback.
 For all baseline models, we set the maximum generation length, sampling temperatures, and other hyperparameters follow the recommended configurations from their respective papers or reports.
@@
 GLM-4.5V (108B) 76.20 67.34 47.00 96.20 87.95 38.60 70.03 69.62
-MiMo-VL-7B-SFT-2508 73.80 68.07 39.00 94.40 91.57 36.40 68.40 67.38
-Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 37.87 67.51 66.55
-Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 66.18 68.64
-FigR (ours) 79.58 79.32 54.00 95.00 93.98 41.92 70.33 73.45
-Gains +6.25 +13.12 +11.00 +1.40 +9.64 +4.05 +2.82 +6.90
+Qwen3-VL-8B-Instruct 61.46 46.20 30.00 94.20 81.93 40.07 67.21 60.15
+Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 41.18 69.73 67.34
+Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 70.33 69.23
+FigR (ours) 79.58 79.32 54.00 95.00 93.98 44.49 72.40 74.11
+Gains +6.25 +13.12 +11.00 +1.40 +9.64 +3.31 +2.67 +6.77
 Table 1: Main results (%) of FigR on seven mathematical reasoning benchmarks compared with several competitive baselines.
@@
 4.2 Main Results
-As shown in Table 1, FigR achieves substantial gains by incorporating visual cues through end-to-end reinforcement learning, reaching an average accuracy of 73.45% across seven mathematical benchmarks, exceeding the base model (Qwen3-VL-32B-Instruct) by 6.90% and the text-only RL baseline by 4.81%.
-Notably, the improvements are particularly pronounced on challenging benchmarks such as AIME 2025 and BeyondAIME.
-FigR significantly outperforms equivalently sized text-only models (Qwen3-32B Thinking) as well as a larger LVLM GLM-4.5V.
-These results indicate that purely textual chain-of-thought reasoning faces clear limitations on hard problems, whereas multimodal chain-of-thought effectively unlocks the reasoning capacity of FigR.
+As shown in Table 1, FigR achieves substantial gains by learning to integrate executable visual construction into multi-turn reasoning.
+Across seven mathematical benchmarks, FigR achieves an average accuracy of 74.11%, exceeding the base policy (Qwen3-VL-32B-Instruct) by 6.77% and the text-only RL baseline by 4.88%.
+Performance gains are particularly pronounced on challenging benchmarks such as AIME 2025 and BeyondAIME, where reasoning requires maintaining global geometric or structural constraints.
+Notably, FigR outperforms both equivalently sized text-only reasoning models (e.g., Qwen3-32B Thinking) and a larger LVLM GLM-4.5V that lacks executable visual construction.
+These results attribute the gains to learning a policy that selectively constructs controllable, executable visual states as intermediate representations during multi-turn reasoning.
 Figure 4: Ablation analysis across training steps.
-We report average response length, code count, code ratio, average code lines per executable block, and code pass rate under different settings.
-FigR enables more frequent and structured visual reasoning while preserving stable execution behavior, compared to ablated settings.
+We track response length, code ratio, code lines, and code pass rate under different settings.
+FigR exhibits sustained and structured executable visual construction with stable execution behavior, whereas removing the adaptive reward mechanism causes code usage to collapse during training.
 4.3 Ablation Study
-As shown in the Table 2, we conduct ablation studies on two representative datasets: AIME 2025 and BeyondAIME, to evaluate the contributions of different components in FigR.
+As shown in Table 2, we conduct ablation studies on two representative datasets: AIME 2025 and BeyondAIME, to evaluate the contributions of different components in FigR.
 Prompt Engineering (PE) on the Base Model.
@@
 While this leads to moderate performance gains on some datasets, the absence of a learning signal prevents the model from internalizing when and how such behaviors should be applied.
-As a result, the induced multimodal reasoning behaviors remain unstable, which inherently limits the attainable performance gains.
+As a result, the induced drawing behaviors remain unstable, which inherently limits the attainable performance gains.
 Model
@@
    + SFT 53.33 38.00
-   + Text-only RL 69.22 46.00
+   + Text-only RL 73.33 46.00
    + Bagel generated img 64.32 41.00
@@
 DeepEyes 2.34 0.00
-FigR (ours) 79.32 52.25
+FigR (ours) 79.32 54.00
    w/o ARM 70.00 49.00
-   w/o img feedback 73.33 47.00
+   w/o visual feedback 76.67 47.00
 Table 2: Ablation Study on FigR. The base model is Qwen3-VL-32B-Instruct, “PE” denotes prompt engineering, “SFT” denotes supervised fine-tuning, and “ARM” denotes the adaptive reward mechanism.
@@
 Injecting Visual Information without Active Visual Reasoning.
-We evaluate two baselines that directly inject visual information(+ Bagel img and + Qwen img), which generate a diagram conditioned on the problem text by using Bagel-7B-MoT and Qwen-Image wu2025qwen, respectively.
-Although both baselines incorporate visual inputs, neither yields consistent improvements over the base model or text-only RL.
-These results indicate that passively injecting images fails to provide reliable structural guidance and may instead introduce noise.
-This contrast highlights the importance of active, feedback-driven visual reasoning in FigR.
+We evaluate two baselines that provide passively generated diagrams as static context (+Bagel img and +Qwen img).
+These diagrams are generated from the problem text using Bagel-7B-MoT and Qwen-Image wu2025qwen, respectively, but are not constructed or revised during reasoning.
+Although these baselines expose the model to visual content, they do not consistently improve performance over the base model or text-only RL baseline.
+These results show that passively injected visual inputs, without executable construction or feedback integration, fail to serve as reliable intermediate states.
+Without a closed reasoning-execution loop, such images neither enforce global constraints nor provide actionable feedback for subsequent reasoning steps.
 Performance of UMMs and TAVLMs.
@@
 Consequently, neither paradigm provides reliable and informative visual feedback for complex mathematical reasoning.
-Due to space limitations, their full results are deferred to Appendix C
+Due to space limitations, their full results are deferred to Appendix C.
 Ablations on FigR Components.
@@
 (2) Removing visual feedback.
-Building on (1), we further remove visual feedback during reinforcement learning.
-The model still performs multi-turn rollouts and executes code, but receives only textual feedback.
-This setting isolates the contribution of visual feedback in forming a stable multimodal reasoning loop.
-Both ablations lead to substantial performance degradation compared to FigR, demonstrating the effectiveness of the proposed components.
+We remove only visual feedback during reinforcement learning.
+The model still performs multi-turn rollouts and executes code, but receives only textual feedback and its corresponding calling reward.
+This setting isolates the contribution of visual feedback in forming a stable reasoning loop.
+Both ablations lead to performance degradation compared to the full FigR, demonstrating the effectiveness of the proposed components.
 Figure 5: Case study of reasoning behaviors.
 FigR demonstrates more effective integration of visual feedback into the reasoning process, resulting in clearer intermediate reasoning and improved final answers, while the baseline models rely primarily on textual reasoning.
-Evolution of active visual thinking during training.
+Evolution of Active Visual Construction During Training.
 We further analyze the emergence of visual-thinking behaviors on the BeyondAIME dataset, using four complementary metrics: response length (measured in tokens with the Qwen3-VL-32B-Instruct tokenizer), code ratio (fraction of samples that invoke code), average code lines (non-empty lines per code block), and code pass rate (fraction of code blocks that execute successfully).
-(1) When visual feedback is removed (w/o vis), the model exhibits persistently low code usage despite retaining textual execution feedback.
-This behavior highlights a critical limitation: textual feedback fails to introduce structurally novel or irreducible information beyond what the model can already generate in natural language.
-Consequently, the model learns to bypass code execution and instead relies on increasingly verbose textual reasoning, as reflected by longer responses but persistently low code ratio.
-(2) Removing the adaptive reward mechanism (w/o ARM) leads to a transient increase in early code usage.
+(1) Removing the adaptive reward mechanism (w/o ARM) leads to a brief increase in early code usage.
 However, this behavior rapidly collapses: both code ratio and code length drop to zero as the model learns that indiscriminate code invocation does not improve outcomes.
@@
 (2) In contrast, FigR maintains a consistently high code ratio throughout training, accompanied by stable and substantially longer code blocks with a high execution pass rate.
-Visual feedback introduces an external representational modality that cannot be easily replicated within the language space, breaking the text-only reasoning loop.
-Consequently, code execution becomes a necessary and reliable mechanism for constructing and validating global structure, rather than an opportunistic or exploratory behavior.
+Executable visual feedback provides an external state that enables explicit validation of global structure, which cannot be reliably enforced through implicit textual representations alone.
+Consequently, visual thinking becomes a reliable mechanism for validating the whole reasoning process, rather than an opportunistic or exploratory behavior.
 4.4 Case Study
@@
 Consequently, reasoning remains largely local and algebraic, and incorrect assumptions are difficult to detect or correct.
-In contrast, FigR closes the reasoning loop by incorporating visual feedback, enabling it to observe global structural patterns and adjust subsequent reasoning steps accordingly.
+In contrast, FigR closes the reasoning loop by incorporating explicit visual state, enabling it to observe global structural patterns and adjust subsequent reasoning steps accordingly.
 This leads to a more robust and interpretable solution.
-Overall, this case study shows that the advantage of FigR does not stem merely from executing intermediate programs, but from actively integrating visual feedback into the reasoning process.
+Overall, this case study shows that the advantage of FigR does not stem merely from executing intermediate programs, but from actively integrating visual construction into the reasoning process.
 Compared with purely textual reasoning or execution with text-only feedback, active visual thinking provides complementary information that substantially improves reasoning reliability.
 5 Conclusion
-We introduce FigR, which integrates active visual thinking into multi-turn reasoning through end-to-end reinforcement learning.
-By closing the loop between reasoning and visual evidence, FigR enables models to reason over global structural properties that are difficult to infer from text alone.
-An adaptive reward mechanism regulates when visual reasoning should be invoked, promoting selective and reliable figure construction without unnecessary generation.
-Extensive experiments highlight the importance of actively incorporating visual feedback during reasoning and suggest a promising direction for interpretable multimodal reasoning.
+We introduce FigR, which integrates executable visual construction into multi-turn reasoning via end-to-end reinforcement learning.
+By embedding controllable and revisable visual states within the reasoning loop, FigR enables reliable reasoning over global structural constraints that are difficult to maintain through text alone.
+Extensive experiments demonstrate that learning when and how to externalize intermediate structure as executable visual states significantly improves reasoning stability and accuracy on complex mathematical problems.
+Limitations
+While our results demonstrate the effectiveness of executable visual construction for complex reasoning, several limitations remain that suggest directions for future work.
+The proposed method incurs additional computational cost due to multi-turn reasoning and code execution.
+This overhead is inherent to approaches that externalize intermediate states, motivating future efforts toward more efficient reasoning–execution loops through compact visual representations or adaptive control strategies.
+Executable visual construction is particularly beneficial for problems that require maintaining global structural or relational constraints.
+In contrast, tasks that can be easily solved through local or purely symbolic reasoning may see less benefit from visual externalization.
+More precise identification and characterization of such task regimes remains an open research question.
+Our evaluation focuses on mathematical reasoning benchmarks that emphasize structural consistency and verifiable outcomes.
+Extending this paradigm to other domains, including scientific reasoning, program synthesis, and long-horizon planning, represents a natural and promising direction for future work.
 Model
@@
 GLM-4.5V (108B) 76.20 67.34 47.00 96.20 87.95 38.60 70.03 69.62
-MiMo-VL-7B-SFT-2508 73.80 68.07 39.00 94.40 91.57 36.40 68.40 67.38
-Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 37.87 67.51 66.55
+Qwen3-VL-8B-Instruct 61.46 46.20 30.00 94.20 81.93 40.07 67.21 60.15
+Qwen3-VL-32B-Instruct 73.33 66.20 43.00 93.60 84.34 41.18 69.73 67.34
 Unified Multimodal Models
@@
 Chain-of-Focus 1.30 0.00 0.00 1.20 0.00 0.74 0.45 0.53
-Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 66.18 68.64
-FigR (ours) 79.58 79.32 54.00 95.00 93.98 41.92 70.33 73.45
-Gains +6.25 +13.12 +11.00 +1.40 +9.64 +4.05 +2.82 +6.90
-Table 3: Full experimental results (%) on seven mathematical reasoning benchmarks. Green-colored font indicates the improvement of FigR over the baseline Qwen3-VL-32B-Instruct.
-Figure 6: Prompt template for the figure-steered suitability classifier.
+Text-only RL 73.33 69.22 46.00 94.40 87.95 43.38 70.33 69.23
+FigR (ours) 79.58 79.32 54.00 95.00 93.98 44.49 72.40 74.11
+Gains +6.25 +13.12 +11.00 +1.40 +9.64 +3.31 +2.67 +6.77
+Table 3: Full experimental results (%) on seven mathematical reasoning benchmarks compared with several competitive baselines.
+Green-colored font indicates improvement over the baseline Qwen3-VL-32B-Instruct.
+Figure 6: Prompt template for the suitability classifier.
 Appendix A Dataset Details
@@
 B.1 Prompt Template for Suitability Classifier
-We provide a dedicated prompt template for the figure-steered suitability classifier as introduced in Section 3.3 and Section 4.1, which is used to determine whether a given problem is suitable for visual-guided reasoning.
+We provide a dedicated prompt template for the suitability classifier as introduced in Section 3.3 and Section 4.1, which is used to determine whether a given problem is suitable for diagrammatic reasoning.
 As shown in Figure 6, the template instructs the model to assess the necessity and usefulness of visual sketches based solely on the problem statement.
@@
 We adopt three different categories of prompt templates corresponding to the reasoning settings.
-(1) As shown in Figure 7, a multi-turn prompt with drawing feedback is used for FigR, the variant without visual feedback, and the variant without ARM.
+(1) As shown in Figure 7, a multi-turn prompt is used for FigR, the variant without visual feedback, and the variant without ARM.
 Models can iteratively generate output and receive feedback from the interpreter.
-(2) As shown in Figure 8, a standard single-turn prompt is adopted for text-only RL and other baselines, requiring the model to directly output the final answer without intermediate interactions.
-(3) For DeepEyes and Chain-of-Focus, we adopt the prompt templates recommended by the original papers.
+(2) As shown in Figure 8, a standard single-turn prompt is adopted for the text-only RL baseline and other models, requiring the model to directly output the final answer without intermediate interactions.
+(3) For DeepEyes and Chain-of-Focus, we adopt the prompt templates recommended by their original papers.
 Figure 7: Prompt template for multi-turn CoT.